# Steering Flexible Linear Objects in Planar Environments by Two Robot Hands Using Euler’s Elastica Solutions

### v3

@@ I Introduction
Robotic manipulation of flexible linear objects such as cables, wires and fresh food items forms a special challenge in robot grasp mechanics. In these problems, one or two robot hands apply endpoint forces and torques that together with external influences such as gravity and contacts with the environment affect the object shape during manipulation. Robotic applications include cable routing and untangling [14, 20, 29], surgical suturing [12, 13], knot tying [31, 30], compliant mechanisms [27], fresh food handling [32], architectural elements fabrication [6] and agricultural robotics [1].
@@
Related work: We use the term flexible cable for flexible linear objects manipulated in two-dimensions. The modeling of these objects relies on Euler-Bernoulli bending moment law [10]. When a flexible cable is held at an equilibrium state by endpoint forces and moments in the plane, the cable curvature at every point is proportional to the cable bending moment at this point. The coefficient of proportionality is the cable stiffness at each point.
@@
The mechanics literature has used the Euler-Bernoulli law to obtain solutions for flexible cable equilibrium shapes in two-dimensions [18]. The solutions, called Euler’s elastica, form a one-parameter family of periodic shapes aligned along a linear axis [17]. The axis is parallel to the opposing forces applied at the cable endpoints and passes through the zero curvature cable points. While flexible cable equilibrium shapes have been fully characterized in the mechanics literature, verification of their stability typically relies on numerical techniques [3, 16]. Sachkov [24, 23, 25] describes analytic bounds for flexible cable stability that are used in this paper’s steering scheme.
@@
In the robotics literature, sampling based approaches are used to plan flexible cable steering paths. Moll [19] samples cable endpoint positions, then computes their stable shapes by numerical optimization of the flexible cable total elastic energy. Bretl [7] uses Sachkov’s approach to describe the cable total elastic energy minimization as an optimal control problem. Bretl shows that the adjoint differential equation which describes the equilibrium shapes of a flexible cable held by robot hands is fully determined by a small number of co-state variables [7]. In 2-D environments these are endpoint forces and moments (three variables) while in 3-D environments these are endpoint forces and toques (six variables).
@@
Using this insight, Bretl developed sampling based planners for steering flexible cables in 2-D and 3-D environments [7]. However, each sampled co-state requires solution of the adjoint differential equation to obtain the cable shape, then computation of the cable co-state to endpoint Jacobian to verify that each sampled cable shape is stable. Sintov [26] extended this work into a two stage approach. First a roadmap of stable cable shapes is computed for sampled co-states. Then a cable steering path is computed in the physical environment using numerical self-intersection and obstacle avoidance tests.
@@
While this paper focuses on Euler’s elastica as a means to model flexible object shapes, robotic manipulation of heterogeneous flexible objects in agricultural robotics [33] and fresh food handling [32] requires complementary approaches. Papers such as [1], [34] and [15] describe adaptive and learning based approaches to manipulation of such complex flexible objects. One hopes that this paper’s modeling approach will become part of adaptive and learning based manipulation techniques for complex flexible objects.
@@
Paper contributions: This paper considers steering flexible cables in two-dimensions using two robot hands. By maintaining equal endpoint tangents, the cable steering problem is reduced to five configuration variables: the cable base-frame configuration and the cable endpoints relative position. The paper starts with a derivation of Euler’s elastica solutions for flexible cable equilibrium shapes under the optimal control approach. The derivation complements Bretle [7] that focused on the adjoint equation and the use of its co-states to steer flexible cables in 3-D environments. This paper focuses on flexible cable steering in two-dimensions in order to take advantage of the closed form elastica solutions.
@@ II Euler’s Elastica as an Optimal Control Problem
where E>0E\!>\!0 is the cable Young’s modulus of elasticity and I>0I\!>\!0 is the cable cross-sectional 2’nd moment of inertia [28]. The cable stiffness, $EI$, is assumed to be a known parameter.
@@
where $\mbox{$\lambda$}(s)\!=\!(\mbox{$\lambda$}_{x}(s),\mbox{$\lambda$}_{y}(s),\mbox{$\lambda$}_{\phi}(s))$ are the co-state variables. The co-states $(\mbox{$\lambda$}_{x}(s),\mbox{$\lambda$}_{y}(s))$ and $\mbox{$\lambda$}_{\phi}(s)$ represent internal force and bending moment that develop at the cable points. The energy extremal cable shapes represent equilibrium states of the cable held by the robot hands. Along an extremal cable shape, the vector $\mbox{$\lambda$}(s)$ is determined by the adjoint equation [21]
@@
where $s\!\in\![0,L]$. When a flexible cable is steered with equal endpoint tangents, its equilibrium shapes possess inflection points, zero curvature points at which the cable curvature switches sign (Fig. 2). Such cable shapes are called inflectional elastica. [^1] Non-inflectional elastica form spiral-like shapes when held with equal endpoint tangents. Such shapes may be useful in robotic applications, but the current paper focuses on the more common inflectional elastica. The curvature of the inflectional elastica is given by an elliptic cosine function of the cable path length parameter [18][p. 402-404]
@@
The elliptic cosine function is a periodic cosine-like function having two zeroes per period [22]. The elastica axis shown in Fig. 2 is defined as the line that passes through the elastica inflection points. Some intuition on this axis can be gained by the following observations. In Eq. (8), the term $\mbox{\small$EI$}\!\cdot\!(\mbox{$\kappa$}\mbox{\small$(s)$}\!-\!\mbox{$\kappa$}(\mbox{\small$0$}))$ multiplies the column vector $(\sin\phi_{0},-\!\cos\phi_{0})$. Since $\mbox{$\kappa$}(s)$ is a periodic cosine-like function, $\phi_{0}$ is the elastica axis angle. It also follows from Eq. (8) that the elastica height above its axis is $h(s)\!=\!\mbox{\small$A$}\!\cdot\!\mbox{\sf$cn$}(\mbox{\small$\sqrt{\mbox{\small$\lambda$}}$}\!\cdot\!(s\!+\!s_{0}),\mathrm{k})$ (Fig. 3), which explains the amplitude parameter $A$.
@@
The elastica full period length: The elliptic cosine and sine functions period is $4K(\mathrm{k})$, where $K(\mathrm{k})$ is the complete elliptic integral of the first kind. [^2] It is defined as $K(\mathrm{k})\!=\!\int_{0}^{\pi/2}\!\!\tfrac{d\theta}{\sqrt{1-\mathrm{k}^{2}\sin^{2}\theta}}$, a standard analytic function of the modulus parameter $\mathrm{k}$ [22]. Denote by $\tilde{L}$ the full period length of the elastica that contains the physical cable. The argument $\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot(s\!+\!s_{0})$ in Eqs. (9)-(10) satisfies the full-period relation $\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot\tilde{L}\!=\!4K(\mathrm{k})$, which gives $\tilde{L}\!=\!4K(\mathrm{k})/\mbox{\small$\sqrt{\mbox{$\lambda$}}$}$.
@@ IV Flexible Cable Stability
This section uses the elastica solutions to obtain a simple geometric characterization of the stable cable shapes to be used during steering. Recall that the elastica inflection points are zero curvature points (Fig. 2). The following theorem adapted from Sachkov [24] characterizes the stable cable shapes in terms of the number of inflection points.
@@
One technique to identify a conjugate point (which marks loss of stability) is described in [9][Definition 27.4]. Consider two families of energy extremal cable shapes sharing a fixed initial point and fixed tangent at this point, parameterized by $\lambda$, $\mathrm{k}$ and $s_{0}$. If the two distinct families merge into a single extremal cable shape at the limit $(\mbox{\small$\lambda$},\mathrm{k},s_{0})\rightarrow(\mbox{\small$\lambda$}^{*},\mathrm{k}^{*},s^{*}_{0})$ while intersecting at their respective endpoints, the endpoint of the limit cable shape associated with $(\mbox{\small$\lambda$}^{*},\mathrm{k}^{*},s^{*}_{0})$ forms a conjugate point along the limit cable shape.
@@ V Flexible Cable Approximation by Quadratic Arcs
Bézier polynomial construction: Let the flexible cable control points be indexed according to their appearance on the cable, $p_{0},\ldots,p_{n}$ such that $p_{0}$ and $p_{n}$ are the cable endpoints and $n\!\leq\!5$ (Fig. 8). Every pair of control points, $p_{i}$ and $p_{i+1}$, bounds a convex arc approximated by a convex quadratic arc as follows. The control points $p_{i}$ and $p_{i+1}$ are first augmented with an intermediate control point, $q_{i}$, located at the intersection point of the cable’s tangent lines at $p_{i}$ and $p_{i+1}$ (Fig. 8). The position of $p_{i}$ and $p_{i+1}$ and the tangents at these points are obtained from the elastica solution of Eq. (8). This data feeds a simple formula for computing $q_{i}$ which is omitted. Using Bézier’s technique [5], each triplet $p_{i}$, $q_{i}$ and $p_{i+1}$ defines the quadratic polynomial
@@ VI Flexible Cable Steering Scheme
    if $L<\tilde{L}$ then set $s_{0}=s_{0,1}$ $s_{0,1}=s_{0,2}$ in this case*/
    else set $s_{0}\!=\!s_{0,1}$ when endpoint curvatures are positive and
    set $s_{0}\!=\!s_{0,2}$ when endpoint curvatures are negative.
    $P(t)\leftarrow BezierPolynomial(\mathrm{k},s_{0},\mbox{\scriptsize$\tilde{L})$}$
    $Q(t)\leftarrow P(t)\,\mathrm{positioned}\,\mathrm{at}\,(x(0),y(0),\phi(0))$.
    if $Collision(Q(t),\mbox{$\cal B$}_{1},\ldots,\mbox{$\cal B$}_{k})=\mbox{\scriptsize$\mathrm{FALSE}$}$ then add $z$ to $\cal O$.
@@ VII Representative Examples
This section describes execution examples of the steering scheme described in Section VI. The scheme was implemented on MATLAB using elliptic functions library [2]. [^3] All simulations run MATLAB-22B on DELL OptiPlex 7080, with Intel I7-10700 CPU running at 2.90 GHz and 16 GB main memory. In all examples the flexible cable maximal flattening parameter is set to $\rho\!=\!0.5$, so that $\tilde{L}$ is at most twice the physical cable length.
@@ Acknowledgment
The derivative of $\mbox{\sf$cn$}(u,\mathrm{k})$ is given by the rule: $\tfrac{d}{du}\mbox{\sf$cn$}(u,\mathrm{k})=-\mbox{\sf$sn$}(u,\mathrm{k})\cdot\mbox{\sf$dn$}(u,k)$, where $\mbox{\sf$sn$}(u,\mathrm{k})$ is the elliptic sine function and $\mbox{\sf$dn$}(u,\mathrm{k})$ is the elliptic delta amplitude [22]. Applying this derivative rule to $\mbox{$\kappa$}(s)$ in Eq. (22) gives
@@
Taking the derivative of both sides of Eq. (29) with respect to $s$ using the derivative rule $\tfrac{d}{du}\mbox{\sf$sn$}(u,\mathrm{k})\!=\!\mbox{\sf$cn$}(u,\mathrm{k})\cdot\mbox{\sf$dn$}(u,k)$ [22] gives
@@
Using the identity $\mbox{\sf$dn$}^{2}(u,\mathrm{k})\!+\!\mathrm{k}^{2}\mbox{\sf$sn$}^{2}(u,\mathrm{k})\!=\!1$ [22], it suffices to verify the relation
@@
We now write the equations of the elastica curve $(x(s),y(s))$ according to the parametrization in [8]:
@@
In [8], there is an expression for the height $y_{fold}$, which corresponds to what we referred to as fold points. From this expression, it is possible to derive an expression for $s_{fold}$, the length along the cable where the tangent angle with respect to the elastica axis satisfies $\phi(s_{fold})\pm 90^{\circ}$:
@@ XI Region 3
Important note: This paper do not consider Region 3 of $\mathrm{k}$-values for local stability reasons. However, locally stable shapes can exist in Region 3. To determine whether a shape is locally stable, one must check the determinant of the endpoint mapping Jacobian [23].

### v4

@@ I Introduction
Robotic manipulation of flexible linear objects such as cables, wires and fresh food items forms a special challenge in robot grasp mechanics. In these problems, one or two robot hands apply endpoint forces and torques that together with external influences such as gravity and contacts with the environment affect the object shape during manipulation. Robotic applications include cable routing and untangling [cable_routing11, bretl_multiple, cable_rss22], surgical suturing [jackson&cavusogl, javdani], knot tying [balkcom15, hirai_knot], compliant mechanisms [till17], fresh food handling [food_survey22], architectural elements fabrication [brander] and agricultural robotics [adaptive_LDO22].
@@
Related work: We use the term flexible cable for flexible linear objects manipulated in two-dimensions. The modeling of these objects relies on Euler-Bernoulli bending moment law [goss_thesis]. When a flexible cable is held at an equilibrium state by endpoint forces and moments in the plane, the cable curvature at every point is proportional to the cable bending moment at this point. The coefficient of proportionality is the cable stiffness at each point.
@@
The mechanics literature has used the Euler-Bernoulli law to obtain solutions for flexible cable equilibrium shapes in two-dimensions [love]. The solutions, called Euler’s elastica, form a one-parameter family of periodic shapes aligned along a linear axis [levyakov10]. The axis is parallel to the opposing forces applied at the cable endpoints and passes through the zero curvature cable points. While flexible cable equilibrium shapes have been fully characterized in the mechanics literature, verification of their stability typically relies on numerical techniques [batista15, levyakov09]. Sachkov [sachkov07, sachkov_conjugate, sachkov_vertices] describes analytic bounds for flexible cable stability that are used in this paper’s steering scheme.
@@
In the robotics literature, sampling based approaches are used to plan flexible cable steering paths. Moll [kavraki06] samples cable endpoint positions, then computes their stable shapes by numerical optimization of the flexible cable total elastic energy. Bretl [bretl_ijrr14] uses Sachkov’s approach to describe the cable total elastic energy minimization as an optimal control problem. Bretl shows that the adjoint differential equation which describes the equilibrium shapes of a flexible cable held by robot hands is fully determined by a small number of co-state variables [bretl_ijrr14]. In 2-D environments these are endpoint forces and moments (three variables) while in 3-D environments these are endpoint forces and toques (six variables).
@@
Using this insight, Bretl developed sampling based planners for steering flexible cables in 2-D and 3-D environments [bretl_ijrr14]. However, each sampled co-state requires solution of the adjoint differential equation to obtain the cable shape, then computation of the cable co-state to endpoint Jacobian to verify that each sampled cable shape is stable. Sintov [sintov20] extended this work into a two stage approach. First a roadmap of stable cable shapes is computed for sampled co-states. Then a cable steering path is computed in the physical environment using numerical self-intersection and obstacle avoidance tests.
@@
While this paper focuses on Euler’s elastica as a means to model flexible object shapes, robotic manipulation of heterogeneous flexible objects in agricultural robotics [zhu_survey22] and fresh food handling [food_survey22] requires complementary approaches. Papers such as [adaptive_LDO22], [zhu_iros18] and [lagneau20] describe adaptive and learning based approaches to manipulation of such complex flexible objects. One hopes that this paper’s modeling approach will become part of adaptive and learning based manipulation techniques for complex flexible objects.
@@
Paper contributions: This paper considers steering flexible cables in two-dimensions using two robot hands. By maintaining equal endpoint tangents, the cable steering problem is reduced to five configuration variables: the cable base-frame configuration and the cable endpoints relative position. The paper starts with a derivation of Euler’s elastica solutions for flexible cable equilibrium shapes under the optimal control approach. The derivation complements Bretle [bretl_ijrr14] that focused on the adjoint equation and the use of its co-states to steer flexible cables in 3-D environments. This paper focuses on flexible cable steering in two-dimensions in order to take advantage of the closed form elastica solutions.
@@ II Euler’s Elastica as an Optimal Control Problem
where E>0E\!>\!0 is the cable Young’s modulus of elasticity and I>0I\!>\!0 is the cable cross-sectional 2’nd moment of inertia [timoshenko&goodier:1969]. The cable stiffness, $EI$, is assumed to be a known parameter.
@@
where $\mbox{$\lambda$}(s)\!=\!(\mbox{$\lambda$}_{x}(s),\mbox{$\lambda$}_{y}(s),\mbox{$\lambda$}_{\phi}(s))$ are the co-state variables. The co-states $(\mbox{$\lambda$}_{x}(s),\mbox{$\lambda$}_{y}(s))$ and $\mbox{$\lambda$}_{\phi}(s)$ represent internal force and bending moment that develop at the cable points. The energy extremal cable shapes represent equilibrium states of the cable held by the robot hands. Along an extremal cable shape, the vector $\mbox{$\lambda$}(s)$ is determined by the adjoint equation [pontryagin]
@@
where $s\!\in\![0,L]$. When a flexible cable is steered with equal endpoint tangents, its equilibrium shapes possess inflection points, zero curvature points at which the cable curvature switches sign (Fig. 2). Such cable shapes are called inflectional elastica. [^1] Non-inflectional elastica form spiral-like shapes when held with equal endpoint tangents. Such shapes may be useful in robotic applications, but the current paper focuses on the more common inflectional elastica. The curvature of the inflectional elastica is given by an elliptic cosine function of the cable path length parameter [love][p. 402-404]
@@
The elliptic cosine function is a periodic cosine-like function having two zeroes per period [ellip_book]. The elastica axis shown in Fig. 2 is defined as the line that passes through the elastica inflection points. Some intuition on this axis can be gained by the following observations. In Eq. (8), the term $\mbox{\small$EI$}\!\cdot\!(\mbox{$\kappa$}\mbox{\small$(s)$}\!-\!\mbox{$\kappa$}(\mbox{\small$0$}))$ multiplies the column vector $(\sin\phi_{0},-\!\cos\phi_{0})$. Since $\mbox{$\kappa$}(s)$ is a periodic cosine-like function, $\phi_{0}$ is the elastica axis angle. It also follows from Eq. (8) that the elastica height above its axis is $h(s)\!=\!\mbox{\small$A$}\!\cdot\!\mbox{\sf$cn$}(\mbox{\small$\sqrt{\mbox{\small$\lambda$}}$}\!\cdot\!(s\!+\!s_{0}),\mathrm{k})$ (Fig. 3), which explains the amplitude parameter $A$.
@@
The elastica full period length: The elliptic cosine and sine functions period is $4K(\mathrm{k})$, where $K(\mathrm{k})$ is the complete elliptic integral of the first kind. [^2] It is defined as $K(\mathrm{k})\!=\!\int_{0}^{\pi/2}\!\!\tfrac{d\theta}{\sqrt{1-\mathrm{k}^{2}\sin^{2}\theta}}$, a standard analytic function of the modulus parameter $\mathrm{k}$ [ellip_book]. Denote by $\tilde{L}$ the full period length of the elastica that contains the physical cable. The argument $\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot(s\!+\!s_{0})$ in Eqs. (9)-(10) satisfies the full-period relation $\mbox{\small$\sqrt{\mbox{$\lambda$}}$}\cdot\tilde{L}\!=\!4K(\mathrm{k})$, which gives $\tilde{L}\!=\!4K(\mathrm{k})/\mbox{\small$\sqrt{\mbox{$\lambda$}}$}$.
@@ IV Flexible Cable Stability
This section uses the elastica solutions to obtain a simple geometric characterization of the stable cable shapes to be used during steering. Recall that the elastica inflection points are zero curvature points (Fig. 2). The following theorem adapted from Sachkov [sachkov07] characterizes the stable cable shapes in terms of the number of inflection points.
@@
One technique to identify a conjugate point (which marks loss of stability) is described in [gelfand][Definition 27.4]. Consider two families of energy extremal cable shapes sharing a fixed initial point and fixed tangent at this point, parameterized by $\lambda$, $\mathrm{k}$ and $s_{0}$. If the two distinct families merge into a single extremal cable shape at the limit $(\mbox{\small$\lambda$},\mathrm{k},s_{0})\rightarrow(\mbox{\small$\lambda$}^{*},\mathrm{k}^{*},s^{*}_{0})$ while intersecting at their respective endpoints, the endpoint of the limit cable shape associated with $(\mbox{\small$\lambda$}^{*},\mathrm{k}^{*},s^{*}_{0})$ forms a conjugate point along the limit cable shape.
@@ V Flexible Cable Approximation by Quadratic Arcs
Bézier polynomial construction: Let the flexible cable control points be indexed according to their appearance on the cable, $p_{0},\ldots,p_{n}$ such that $p_{0}$ and $p_{n}$ are the cable endpoints and $n\!\leq\!5$ (Fig. 8). Every pair of control points, $p_{i}$ and $p_{i+1}$, bounds a convex arc approximated by a convex quadratic arc as follows. The control points $p_{i}$ and $p_{i+1}$ are first augmented with an intermediate control point, $q_{i}$, located at the intersection point of the cable’s tangent lines at $p_{i}$ and $p_{i+1}$ (Fig. 8). The position of $p_{i}$ and $p_{i+1}$ and the tangents at these points are obtained from the elastica solution of Eq. (8). This data feeds a simple formula for computing $q_{i}$ which is omitted. Using Bézier’s technique [bezier], each triplet $p_{i}$, $q_{i}$ and $p_{i+1}$ defines the quadratic polynomial
@@ VI Flexible Cable Steering Scheme
    if $L=\tilde{L}$ then set $s_{0}=s_{0,1}$ $s_{0,1}=s_{0,2}$ in this case*/
    else set $s_{0}\!=\!s_{0,1}$ when $\kappa(0)>0$ and
        set $s_{0}=s_{0,2}$ when $\kappa(0)<0$.
[…]
    $Q(t)\leftarrow\mathrm{Position}\big(P(t),x(0),y(0),\phi(0)\big)$
    if $Collision(Q(t),\mathcal{B}_{1},\ldots,\mathcal{B}_{k})=\mathrm{FALSE}$
    then add $z$ to $\mathcal{O}$
  end for
@@ VII Representative Examples
This section describes execution examples of the steering scheme described in Section VI. The scheme was implemented on MATLAB using elliptic functions library [batista_matlab]. [^3] All simulations run MATLAB-22B on DELL OptiPlex 7080, with Intel I7-10700 CPU running at 2.90 GHz and 16 GB main memory. In all examples the flexible cable maximal flattening parameter is set to $\rho\!=\!0.5$, so that $\tilde{L}$ is at most twice the physical cable length.
@@ Acknowledgment
The derivative of $\mbox{\sf$cn$}(u,\mathrm{k})$ is given by the rule: $\tfrac{d}{du}\mbox{\sf$cn$}(u,\mathrm{k})=-\mbox{\sf$sn$}(u,\mathrm{k})\cdot\mbox{\sf$dn$}(u,k)$, where $\mbox{\sf$sn$}(u,\mathrm{k})$ is the elliptic sine function and $\mbox{\sf$dn$}(u,\mathrm{k})$ is the elliptic delta amplitude [ellip_book]. Applying this derivative rule to $\mbox{$\kappa$}(s)$ in Eq. (22) gives
@@
Taking the derivative of both sides of Eq. (29) with respect to $s$ using the derivative rule $\tfrac{d}{du}\mbox{\sf$sn$}(u,\mathrm{k})\!=\!\mbox{\sf$cn$}(u,\mathrm{k})\cdot\mbox{\sf$dn$}(u,k)$ [ellip_book] gives
@@
Using the identity $\mbox{\sf$dn$}^{2}(u,\mathrm{k})\!+\!\mathrm{k}^{2}\mbox{\sf$sn$}^{2}(u,\mathrm{k})\!=\!1$ [ellip_book], it suffices to verify the relation
@@
We now write the equations of the elastica curve $(x(s),y(s))$ according to the parametrization in [djondjorov]:
@@
In [djondjorov], there is an expression for the height $y_{fold}$, which corresponds to what we referred to as fold points. From this expression, it is possible to derive an expression for $s_{fold}$, the length along the cable where the tangent angle with respect to the elastica axis satisfies $\phi(s_{fold})\pm 90^{\circ}$:
@@ XI Region 3
Important note: This paper do not consider Region 3 of $\mathrm{k}$-values for local stability reasons. However, locally stable shapes can exist in Region 3. To determine whether a shape is locally stable, one must check the determinant of the endpoint mapping Jacobian [sachkov_conjugate].
@@ XIII-B Measurement Errors
A preliminary error measurement was performed on a fully stretched zip-tie. The average markers measurement error was determined to be a disc of radius 1.2 mm.